# Characterization of a fiber laser hydrophone for acoustic neutrino detection

### v1

@@ 4.3.1 Impact of protective cap
As shown in figure 1 a cap is glued on both ends of the transducer to guide the fiber through the membrane of the transducer. The cap was added to the transducer for practical reasons, to avoid breaking the optical fiber during a extensive set of measurements and handling. To determine the impact of the cap on the acoustic characteristics of the hydrophone, the transfer function of the hydrophone has been determined with and without the cap. In figure 10 (a) the transfer function of the single membrane sensor is shown for both configurations. As can be seen from the figure, the cap has a non-negligible impact on the acoustic characteristics: the main mechanical resonance has shifted to a lower eigenfrequency. In addition, a second resonance peak can be seen that is absent when no cap has been applied. The main mechanical resonance peak shifts from 13.9 kHz to a frequency of 12.0 kHz. Furthermore there are two small anti-resonance peaks, at 4.8 and 5.7 kHz respectively. The latter is introduced by the cap and shows directionality, while the first anti-resonance peak is always present, irrespective of the cap or direction w.r.t to the sound source. The origin of the peak is unknown. Note further that the Helmholtz resonance peak remained at its position.
Figure 10: (a) The hydrophone response as a function of the acoustic frequency for a single membrane transducer with a cap compared to a single membrane transducer without a protective cap. In (b) the transfer functions of a sensor with a sensor volume completely filled with water and with an air bubble inserted.
@@ 4.3.2 Air bubbles in the sensor
As mentioned in section 2.0.1, a static pressure mechanism has been implemented by introducing two orifices in the sensor. Prior deployment the sensor volume has been filled with water, which allows to withstand the ambient pressure when the sensor is required to operate at large depth. Filling the sensor volume with water however, (instead of air) does compromise the sensitivity of the sensor. It is important to completely fill the sensor as any residual air will provide a risk to deform and damage the sensor. It is anticipated that residual air is expected to have an impact on the transfer function. To verify the impact of air in the sensor, we have determined the transfer function without air, after careful filling of the sensor and with air. Then, a small bubble of air was inserted in the sensor using an injection needle. Unfortunately we did not have full control of the size of air bubble, but we estimated its size to be small and around 1% percent in volume only with respect to the sensor volume.
In figure 10 (b) we show the transfer function with and without air bubble in the sensor. Comparing the transfer functions differ both in the position of the mechanical resonant frequency as well as the overall strain sensitivity. After introducing air in the sensor volume, the peak shifts to lower frequencies and the overall sensitivity seem to increase. The shape of the transfer function did not change. The impact of the air bubble is confirmed by simulations.
@@ 4.4 Linearity
To check the linearity of the response of the hydrophone the response was determined to a tone of 5 kHz at several setting of the sound source. The sound pressure level of the source was set by a voltage and the recorded pressure level was determined for a single tone of 5 kHz. In figure 11 the recorded height of the tone as a function of the voltage applied to the source is shown. The sensor shows an excellent linear behaviour in the applicable range.
@@ 4.5 Directionality
The experimental set-up allowed to change the orientation of the fiber hydrophone with respect to the source, so that the directionality could be determined. To do so, the transfer function has been determined at various functions for both types of hydrophones. In figure 12 we show the transfer function for various angles with respect to the source for a double membrane sensor. As can be seen from the figure, the transfer function for the double membrane hydrophone varies with the angle, while for the single membrane hydrophone the effect is less pronounced. We furthermore observe that the direction sensitivity is not observed outside the resonance peak. The position of the resonance peak is independent of the angle of incidence.
Figure 12: Directionality of the response. In (a) the transfer function for a double sensor membrane is shown for various angles of the sensor w.r.t. the source. In (b) the resonance peak (normalized to the 90o response) is shown for two types of sensors as a function of the incident angle.
@@ 4.6 Pulse measurement
The investigation of the (complex) transfer function of the fiber hydrophone is justified when we have a look at the single pulses or transients. A number of pulse trains have been recorded to this end. In figure 13 we show an example of one such pulse, measured with both the fiber hydrophone. The impact of the resonant peak is apparent as the signal as measured using the fiber hydrophone shows a large amount of ringing. Knowing the transfer function of the hydrophone, we can deconvolute the transfer function as compare the resulting pulse with the reference hydrophone. To do so we take the transfer function as shown in figure 8 and do interpolation to match the transfer function to the amount of data points in the pulse. The result of the deconvolution is shown in the insert of figure 13. After deconvultion, both pulses are nearly indistinguishable, an indication that the transfer function function has been properly determined. Note that also the reference hydrophone shows some ringing. As the reference hydrophone has a flat frequency response, the ringing is assumed to originate from the source.
Figure 13: Measurement of a single pulse using both the fiber hydrophone and the reference hydrophone. In addition we show the deconvoluted pulse as measured using the fiber hydrophone.
@@ 4.7 Pressure stress tests
One of the challenging requirements to meet is the fact that the hydrophones is operated in the high pressure environments of the deep sea. A first investigation to verify the compliance with the requirement has been carried out in a pressure stress test. To this end a hydrophone was placed in a small hyperbaric pressure vessel and put under pressure of 140 bar. The pressure was applied at several steps. In figure 14 the time line of the measurement is shown: during a time period of 600 seconds the phase measurement in the interferometer has been monitored, while the ambient pressure increased by steps of roughly 20 bar. The pressure was applied manually (pump HP32-8 by Uraca) in steps of roughly 20 bar, while each step was followed by a few seconds of stable pressure until a maximum of 140 bar was reached. Then, the pressure was held to a constant level for about 3 minutes after which the pressure was released and returned to ambient pressure. The graph in figure 14 shows the wavelength shift during the complete pressure cycle. Note that the maximum pressure applied was limited by the fact that the vessel was put under pressure by hand.
The high pressure will slightly deform the transducer and the integrated fiber laser. This gives rise to a phase shift as measured in the interferometer in the interrogator and is plotted in figure 14. Subsequently the wavelength shift of the fiber laser has been determined using equation 5. The figure shows that due to the deformation of the transducer the shift of the central wavelength of the fiber laser will shift 60 pm at a pressure of 140 bar. Once the hydrophone will be deployed to the deep sea, the central wavelength of the laser will remain at this value. Note that this measurement indicates that the pressure compensation does not remove the dependency of the wavelength of the hydrophone on the applied depth completely, but enough for proper operation.
@@
If we zoom in on the time trace of the measurement of the phase (and hence the wavelength) during the pressure qualification as shown in figure 14, then we can plot again the spectrogram. The spectrogram of the period between 354 and 357 seconds after the start of the measurement is shown in the figure 15. In the top panel we show the spectrogram, while in the bottom panel the time trace is plotted. The time trace is plotted after a high-pass filter (with a cut-off frequency of 500 Hz). The time trace shows spiky behaviour: The measured wavelength does not only show a gradual decrease and subsequent increase during the measurement (as shown in figure 14), but additional transient signals have been recorded. The origin of the signal is not clear; this could well be the deformation of the transducer, but also acoustic signals from the pressure vessel itself. We can use these transients to investigate one of the main features of the transfunction, i.e. the resonance peak. The spectrogram shows the position of the resonance peak at $\sim$16 kHz by the dark blue horizontal band. We can therefore conclude that the position of the resonance peak did not alter when the transducer was put onder a static pressure of 140 bar.
@@
A measurement of the transfer function was carried out before and after the pressure cycle in the anechoic basin of which the results. There was no noticeable difference found between the two transfer functions: The mechanical resonance peak was found at the same frequency and the overall strain sensitivity remained identical. This indicates that the sensor and the attachment of the fiber to the transducer did not deform. In addition, the same directionality has been found as depicted in figure 12.
It should be noted that the pressure cycle in this experiment was much shorter than what can be expected during deployment in the applicable environment. In case the hydrophones would be deployed in the deep sea, the rate of pressure change would be an order of magnitude smaller, i.e. the hydrophone would experience a more benign change of environment.
5 Sensitivity
Based on the measured transfer function and system noise, it is possible to determine the sensitivity of the fiber laser hydrophone. To do so we take the strain sensitivities as shown in figure 9 of the two ypes of transducers and the system noise as depicted in figure 6. The resulting sensitivity of the hydrophone as a function of the acoustic frequency is plotted in figure 16. It is illustrative to compare the sensitivity with the noise levels as expected in the deep sea. It is common to express the deep sea noise in terms of a sea state noise ([9]), which runs from sea state 0 (SS0) to SS9. The expression of the sea states is provided in [10, 11] and is given by:
@@ 5 Sensitivity
Some levels have been depicted in figure 16 as well and it can be concluded that the overall sensitivity compares to the lowest expected noise levels in the deep sea, varying between sea state 0 and 1. In particular, the sensitivity is enhanced down below sea state 0 at the position of the resonance peaks. Note further that the sensivities of the two type of hydrophones are comparable away from the resonance peaks. Furthermore, although not shown in the figure, the response of the hydrophones is non-zero well above 20 kHz, which allows for higher frequency signals to be detected.
Figure 16: Sensitivity of the two tested types of fiber hydrophones as a function of frequency. In addition ambient sea state noise levels are drawn for comparison. See text for details on sea state noise.
@@ 6 Conclusions
This paper presents the development of a highly sensitive hydrophone sensor utilizing fiber laser technology, designed for potential use in future acoustic neutrino telescopes. To meet the demands of deep-sea operation, the sensor must function at ambient pressures of at least 10 MPa while detecting transient pressure signals in the range of several mPa. A pressure compensation mechanism has been implemented to offset static pressure levels.
@@
The hydrophone system has been subjected to a detailed investigation of the acoustic performance. A dedicated set up in an anechoid pool has been made for this purpose. Based on the presented results we arrive at the following conclusions. The transfer function of the hydrophone follows the predictions of the finite element analysis that were carried out during the design phase: Two resonance peaks are found in the response of the hydrophone: i) one Helmholtz resonance at lower frequencies ($<$ 500 Hz) and ii) one mechanical resonance of the transducer well above 10 kHz. The exact values of the eigenfrequencies differ from the predicted values, but the dynamical range of the sensors is found to be 200 Hz to 20 kHz, compliant with the main frequency components of a transient neutrino signal.
@@
There is significant difference between the transfer functions of a single and a double membrane type hydrophone. While the double membrane type shows an overall higher sensitivity, it also exhibits a substantial directionality, in particular around the resonance peaks. We converted the measured wavelength shift pressure sensitivity for a direct comparison with ambient sea state noise. It was found that the sensitivity of the technology is between sea state 0 and 1 for most of the dynamic range, which allows for detecting weak acoustic signals generated by neutrino interactions in the deep sea.
@@
Measurements on pulses show the impact of a complex transfer function. As expected, transient signals exhibit significant ringing. However, after deconvolution of the fiber hydrophone response the pulse shape matches the reference pulse, confirming that the transfer function is well-characterized and can be corrected. As the pulse shape and size is directly related to the energy of the interacting neutrino, deconvolution is important. Various approaches can be employed, including the deconvolution method presented here, system identification techniques, or even convolution of a template bank with the hydrophone response in the context of matched filtering. The discussion on how the deconvolution exactly is carried out once the hydrophones are implemented in a full size neutrino telescope is beyond the scope of this paper.
@@
Pressure qualifaction has been carried out to confirm the working principle of the static pressure compensation mechanism. The hydrophone was put in a vessel wherafter a pressure of 140 bar (14 MPa) was applied. The hydrophone showed no degradation in sensitivity after the pressure cycle. Transient signals that were observed during the pressure qualification showed that the mechanical resonance of the transducer did not change prior, during and post qualification. This may lead to the conclusion that the transfer function of the transducer under pressure is similar to the one outside the pressure vessel. Whether the overall sensitivity changes under pressure needs verification in a dedicated experiment. It requires an acoustic set-up in a high pressure environment.
Finally, we have investigated some aspects that are related to the integration of the transducer that might have an impact on the transfer function. The impact of a protective cap on the acoustic performance of the transducer has been investigated and it has been shown that the mechanical eigenresonance is affected. Then, the effect on residual air has a substantial impact on the performance of the transducer by shifting the resonance frequency to lower values and increasing the sensitivity at lower frequencies. It should be noted that when the hydrophone is operated in deep sea, all airbubbles will be squeezed out due to the ambient pressure.
@@
The concept of distributed sensing using optical fibers combined with the results of the acoustic measurements and the pressure qualifcation makes the technology is well suited for the implementation in a large acoustic neutrino telescope.
7 Acknowledgement

### v2

@@ 4.3.1 Impact of protective cap
As shown in figure 1 a cap is glued on both ends of the transducer to guide the fiber through the membrane of the transducer. The cap was added to the transducer for practical reasons, to avoid breaking the optical fiber during a extensive set of measurements and handling. To determine the impact of the cap on the acoustic characteristics of the fiber hydrophone, the transfer function of the fiber hydrophone has been determined with and without the cap. In figure 10 (a) the transfer function of the single membrane sensor is shown for both configurations. As can be seen from the figure, the cap has a non-negligible impact on the acoustic characteristics: the main mechanical resonance has shifted to a lower eigenfrequency. In addition, a second resonance peak can be seen that is absent when no cap has been applied. The main mechanical resonance peak shifts from 13.9 kHz to a frequency of 12.0 kHz. Furthermore there are two small anti-resonance peaks, at 4.8 and 5.7 kHz respectively. The latter is introduced by the cap and shows directionality, while the first anti-resonance peak is always present, irrespective of the cap or direction w.r.t to the sound source. The origin of the peak is unknown. Note further that the Helmholtz resonance peak remained at its position.
Figure 10: (a) The fiber hydrophone response as a function of the acoustic frequency for a single membrane transducer with a cap compared to a single membrane transducer without a protective cap. In (b) the transfer functions of a sensor with a sensor volume completely filled with water and with an air bubble inserted.
@@ 4.3.2 Air bubbles in the sensor
As mentioned in section 2.1, a static pressure mechanism has been implemented by introducing two orifices in the sensor. Prior deployment the sensor volume has been filled with water, which allows to withstand the ambient pressure when the sensor is required to operate at large depth. Filling the sensor volume with water however, (instead of air) does compromise the sensitivity of the sensor. It is important to completely fill the sensor as any residual air will provide a risk to deform and damage the sensor. It is anticipated that residual air is expected to have an impact on the transfer function. To verify the impact of air in the sensor, we have determined the transfer function without air, after careful filling of the sensor and with air. Then, a small bubble of air was inserted in the sensor using an injection needle. Unfortunately we did not have full control of the size of air bubble, but we estimated its size to be small and around 1% percent in volume only with respect to the sensor volume.
In figure 10 (b) we show the transfer function with and without air bubble in the sensor. Comparing the transfer functions differ both in the position of the mechanical resonant frequency as well as the overall strain gauge factor. After introducing air in the sensor volume, the peak shifts to lower frequencies and the overall sensitivity seems to increase. The shape of the transfer function did not change. The impact of the air bubble is confirmed by simulations.
@@ 4.4 Linearity
To check the linearity of the fiber hydrophone, its response to a tone of 5 kHz was determined at various sound source settings. The sound pressure level of the source was set by a voltage and the recorded sound pressure level was determined for a tone of 5 kHz. In figure 11 the recorded height of the sound pressure as a function of the voltage applied to the source is shown. The sensor shows a linear behavior in the applicable range.
@@ 4.5 Directionality
The experimental set-up allowed to change the orientation of the fiber hydrophone with respect to the source, so that the directionality could be determined. To do so, the transfer function has been determined at various orientations relative to the source for both types of fiber hydrophones. In figure 12 we show the transfer function for various angles with respect to the source for a double membrane sensor. As can be seen from the figure, the transfer function for the double membrane fiber hydrophone varies with the angle, while for the single membrane fiber hydrophone the effect is less pronounced. We furthermore observe that the direction sensitivity is not observed outside the resonance peak. The position of the resonance peak is independent of the angle of incidence.
Figure 12: Directionality of the response. In (a) the transfer function for a double sensor membrane is shown for various angles of the sensor with respect to the source. In (b) the resonance peak (normalized to the 90o response) is shown for two types of sensors as a function of the incident angle.
@@ 4.6 Pulse measurement
The investigation of the (complex) transfer function of the fiber hydrophone is justified when we have a look at the single pulses or transients. A number of pulse trains have been recorded to this end. In figure 13 we show an example of one such pulse, measured with both the fiber hydrophone and the reference hydrophone.
The mechanical eigenfrequency of the transducer is manifest as a resonant peak in the transfer function and as a significant amount of ringing in the time trace: a short transient signal shows a large tail that oscillates with the eigenfrequency. Knowing the transfer function of the hydrophone, we can deconvolute the transfer function to compare the resulting pulse with the reference hydrophone. To do so we take the transfer function as shown in figure 8 and do interpolation to match the transfer function to the amount of data points in the pulse. The result of the deconvolution is shown in the insert of figure 13. After deconvolution, both pulses are nearly indistinguishable, an indication that the transfer function function has been properly determined. Note that also the signal recorded by the reference hydrophone shows some ringing. Since the reference hydrophone has a flat frequency response (and hence no ringing is expected), the ringing is assumed to originate from the source.
Figure 13: Measurement of a single pulse using both the fiber hydrophone and the reference hydrophone. In addition we show the deconvoluted pulse measured by the fiber hydrophone.
@@ 4.7 Pressure stress tests
One of the challenging requirements to meet is the fact that the fiber hydrophones are operated in the high pressure environments of the deep sea. A first investigation to verify the compliance with the requirement has been carried out in a pressure stress test. To this end a fiber hydrophone was placed in a small hyperbaric pressure vessel and put under pressure of 140 bar. The pressure was applied at several steps. In figure 14 the time line of the measurement is shown: during a time period of 600 seconds the phase measurement in the interferometer has been monitored, while the ambient pressure increased by steps of roughly 20 bar. The pressure was applied manually (pump HP32-8 by Uraca) in steps of roughly 20 bar, while each step was followed by a few seconds of stable pressure until a maximum of 140 bar was reached. Then, the pressure was held to a constant level for about 3 minutes after which the pressure was released and returned to ambient pressure. The graph in figure 14 shows the wavelength shift during the complete pressure cycle. Note that the maximum pressure applied was limited by the fact that the vessel was put under pressure by hand.
The high pressure will slightly deform the transducer and the integrated fiber laser. This gives rise to a phase shift as measured in the interferometer in the interrogator and is plotted in figure 14. Subsequently the wavelength shift of the fiber laser has been determined using equation 5. The figure shows that due to the deformation of the transducer the shift of the central wavelength of the fiber laser will shift 60 pm at a pressure of 140 bar. Once the fiber hydrophone will be deployed to the deep sea, the central wavelength of the laser will remain at this value. Note that this measurement indicates that the pressure compensation does not remove the dependency of the wavelength of the fiber hydrophone on the applied depth completely, but enough for proper operation.
@@
If we zoom in on the time trace of the measurement of the phase (and hence the wavelength) during the pressure qualification as shown in figure 14, then we can plot again the spectrogram. The spectrogram of the period between 354 and 357 seconds after the start of the measurement is shown in the figure 15. In the top panel we show the spectrogram, while in the bottom panel the time trace is plotted. The time trace is plotted after a high-pass filter (with a cut-off frequency of 500 Hz). The time trace shows spiky behavior: The measured wavelength does not only show a gradual decrease and subsequent increase during the measurement (as shown in figure 14), but additional transient signals have been recorded. The origin of the signal is not clear; this could well be the deformation of the transducer, but also acoustic signals from the pressure vessel itself. We can use these transients to investigate one of the main features of the transfer function, i.e. the resonance peak. The spectrogram shows the position of the resonance peak at $\sim$16 kHz by the dark blue horizontal band. We can therefore conclude that the position of the resonance peak was not altered when the transducer was put under a static pressure of 140 bar.
@@
A measurement of the transfer function was carried out before and after the pressure cycle in the anechoic basin of which the results. There was no noticeable difference found between the two transfer functions: The mechanical resonance peak was found at the same frequency and the overall strain gauge factor remained identical. This indicates that the sensor and the attachment of the fiber to the transducer did not deform. In addition, the same directionality has been found as depicted in figure 12.
It should be noted that the pressure cycle in this experiment was much shorter than what can be expected during deployment in the applicable environment. If the fiber hydrophones were deployed in the deep sea, the rate of pressure change would be an order of magnitude smaller, resulting in a more gradual environmental transition.
4.8 Sensitivity
Based on the measured transfer function and system noise, it is possible to determine the sensitivity of the fiber hydrophone. We therefore take the strain gauge factor as shown in figure 9 of the two types of transducers and the system noise as depicted in figure 6. The resulting sensitivity of the fiber hydrophone as a function of the acoustic frequency is plotted in figure 16. It is illustrative to compare the sensitivity with the noise levels as expected in the deep sea. It is common to express the deep sea noise in terms of a sea state noise ([15]), which runs from sea state 0 (SS0) to SS9. The expression of the sea states is provided in [16, 17] and is given by:
@@ 4.8 Sensitivity
Some levels have been depicted in figure 16 as well and it can be concluded that the overall sensitivity compares to the lowest expected noise levels in the deep sea, varying between sea state 0 and 1. In particular, the sensitivity is enhanced down below sea state 0 at the position of the resonance peaks. Note further that the sensitivities of the two types of fiber hydrophones are comparable away from the resonance peaks. Furthermore, although not shown in the figure, the response of the fiber hydrophones is non-zero well above 20 kHz, which allows for higher frequency signals to be detected. Finally, as an example, the spectrum of a neutrino was added to the figure. The spectrum of the neutrino signal may vary with energy or distance and orientation with respect to the particle shower in the water. In this case a signal was simulation for a neutrino with energy of $10^{10}$ GeV at a distance of 100 m.
Figure 16: Sensitivity of the two tested types of fiber hydrophones as a function of frequency. In addition ambient sea state noise levels are drawn for comparison as well as an example of a neutrino pulse. See text for details on sea state noise.
@@ 5 Conclusions
This paper presents the development of a highly sensitive fiber hydrophone based on fiber laser technology, designed for potential use in future acoustic neutrino telescopes. To meet the demands of deep-sea operation, the sensor must function at ambient pressures of at least 10 MPa while detecting transient pressure signals in the range of several mPa. A pressure compensation mechanism has been implemented to offset static pressure levels.
@@
The fiber hydrophone system has been subjected to a detailed investigation of the acoustic performance. A dedicated set up in an anechoic pool has been made for this purpose. Based on the presented results we arrive at the following conclusions. The transfer function of the fiber hydrophone follows the predictions of the finite element analysis that were carried out during the design phase: Two resonance peaks are found in the response of the fiber hydrophone: i) one Helmholtz resonance at lower frequencies ($<$ 500 Hz) and ii) one mechanical resonance of the transducer well above 10 kHz. The exact values of the eigenfrequencies differ from the predicted values, but the dynamical range of the sensors is found to be 200 Hz to 20 kHz, compliant with the main frequency components of a transient neutrino signal.
@@
There is significant difference between the transfer functions of a single and a double membrane type fiber hydrophone. While the double membrane type shows an overall higher sensitivity, it also exhibits a substantial directionality, in particular around the resonance peaks. We converted the measured wavelength shift pressure sensitivity for a direct comparison with ambient sea state noise. It was found that the sensitivity of the technology is between sea state 0 and 1 for most of the dynamic range, which allows for detecting weak acoustic signals generated by neutrino interactions in the deep sea.
@@
Measurements on pulses show the impact of a complex transfer function. As expected, transient signals exhibit significant ringing. However, after deconvolution of the fiber hydrophone response the pulse shape matches the reference pulse, confirming that the transfer function is well-characterized and can be corrected. As the pulse shape and size is directly related to the energy of the interacting neutrino, deconvolution is important. Various approaches can be employed, including the deconvolution method presented here, system identification techniques, or even convolution of a template bank with the fiber hydrophone response in the context of matched filtering. The discussion on how the deconvolution exactly is carried out once the fiber hydrophones are implemented in a full size neutrino telescope is beyond the scope of this paper.
@@
Pressure qualification has been carried out to confirm the working principle of the static pressure compensation mechanism. The fiber hydrophone was placed in a vessel, after which a pressure of 140 bar (14 MPa) was applied. The fiber hydrophone showed no degradation in sensitivity after the pressure cycle. Transient signals that were observed during the pressure qualification showed that the mechanical resonance of the transducer did not change prior, during and post qualification. This may lead to the conclusion that the transfer function of the transducer under pressure is similar to the one outside the pressure vessel. Whether the overall sensitivity changes under pressure needs verification in a dedicated experiment. It requires an acoustic set-up in a high pressure environment.
@@
The concept of distributed sensing using optical fibers, combined with the results of the acoustic measurements and pressure qualification, makes this technology well-suited for implementation in a large acoustic neutrino telescope.
6 Acknowledgment